OpenStudy (anonymous):

summation from k=1 to 3 (-1)^k (k-2)^2 write the sum without the sigma notation and evaluate it. Help any1?

OpenStudy (anonymous):

\[\sum_{k=1}^{3} (-1)^{k}(k-2)^{2}\]

OpenStudy (loser66):

where are you stuck?

OpenStudy (anonymous):

\[(-1)^{1}(1-2)^{2} + (-1)^{3}(3+2)^{3}\]

OpenStudy (anonymous):

ive tried pluggin in the 1 and the 3

OpenStudy (anonymous):

dunno if im right

OpenStudy (loser66):

where is k =2?

OpenStudy (loser66):

ok, k =1, then the expression is \( (-1)^1(1-2)^2= -1\), right?

OpenStudy (anonymous):

so i have to plug in all the numbers from 1 to 3? lol thought its just 1 and 3

OpenStudy (loser66):

k =2 , then the expression is \((-1)^2 (2-2)^2 =0, right?

OpenStudy (anonymous):

should be from 1 to 3

OpenStudy (loser66):

Yes, although k =2 makes the expression =0, you must put it in to show that it is there but =0

OpenStudy (loser66):

but you calculate k =3 is wrong.

OpenStudy (anonymous):

how?

OpenStudy (loser66):

it is \((-1)^3 (3-2)^2=-1\)

OpenStudy (loser66):

you let (3+2) ^2 not 3-2

OpenStudy (anonymous):

ahaha ye

OpenStudy (anonymous):

so all of that equals to = -2

OpenStudy (loser66):

yup

OpenStudy (anonymous):

thank you very much, have a nice day